# GridShow: Omni Visual Generation

### v3

@@ Abstract
In this paper, we introduce GRID, a novel paradigm that reframes a broad range of visual generation tasks as the problem of arranging grids, akin to film strips. At its core, GRID transforms temporal sequences into grid layouts, enabling image generation models to process visual sequences holistically. To achieve both layout consistency and motion coherence, we develop a parallel flow-matching training strategy that combines layout matching and temporal losses, guided by a coarse-to-fine schedule that evolves from basic layouts to precise motion control. Our approach demonstrates remarkable efficiency, achieving up to 35$\times$ faster inference speeds while using $<\frac{1}{1000}$ of the computational resources compared to specialized models. Extensive experiments show that GRID exhibits exceptional versatility across diverse visual generation tasks, from Text-to-Video to 3D Editing, while maintaining its foundational image generation capabilities. This dual strength in both expanded applications and preserved core competencies establishes GRID as an efficient and versatile omni-solution for visual generation. Our code is available at: https://github.com/Should-AI-Lab/GRID.
Visual Grid Generation
00footnotetext: $\dagger$ Project Lead
Figure 1: Framework Overview: We transforms temporal and view sequences into structured layout spaces, enabling consistent synthesis for diverse generation and interpolation tasks from text and image/frame inputs.
@@ 1 Introduction
Film strips demonstrate an elegant approach in visual arts: by arranging temporal sequences into structured grids, allowing time-based narratives to be displayed in layouts while maintaining their narrative coherence and visual connections. This organization does more than preserve chronological order - it enables efficient content manipulation, comparison, and editing. Drawing inspiration from this intuitive yet powerful organizational principle, we propose a fundamental question: Can we directly reframe various visual generation tasks as how to layout, where key visual elements (such as multiple viewpoints or video frames) are treated as grid “layout”?
To answer this, a natural starting point emerges from the recent breakthroughs in text-to-image generation. Models like  (Esser et al., 2024; Baldridge et al., 2024; Betker et al., 2023) have shown remarkable capabilities in understanding and generating complex spatial relationships. This leads us to a straightforward initial attempt: we test their ability to generate grid-arranged multi-view layouts through simple prompting (Figure 8). However, our experiments reveal that the current models, despite their advanced capabilities, fall short in two fundamental aspects (detailed in Section A.1):
@@
Figure 2: Pipeline Overview. Left: GRID arranges videos into grid layouts, with text annotations describing frame sequences and multi-view spatial relationships. Right: Grid-based reformulation utilizing model’s built-in self-attention.
@@ 2.1 Grid Representation
Our approach demonstrates that thoughtful problem restructuring can be more effective than architectural modifications. By reorganizing the input space into a grid representation, standard text-to-image models can naturally handle temporal generation without architectural changes (see Appendix A.2 for detailed attention mechanism analysis). This grid-based design offers two key advantages: First, it enables parallel generation of all frames and eliminates the error accumulation problems common in autoregressive approaches (Tian et al., 2024). Second, by leveraging the inherent consistency priors within pretrained image generation models, our approach effectively transfers their learned spatial consistency to temporal and multi-view coherence. This crucial advantage avoids the need for extensive pretraining on massive video datasets, as the grid representation naturally extends existing image-level understanding to sequence generation. Additionally, through flexible layout conditioning ($c_{\text{layout}}$), our model shows strong generalization capabilities beyond training constraints (Section A.5), suggesting a promising solution to the fixed-length limitations of existing methods. Additionally, our grid representation supports diverse input types, including multi-view images and multi-frame sequences, laying the foundation for a comprehensive omni-generation model that bridges image and video domains.
@@ 2.2 Parallel Flow Matching
To fully leverage the potential of our grid representation, we employ parallel flow matching (Esser et al., 2024) to ensure temporal coherence across consecutive grids. For each training sample $\mathbf{I}=(I_{ij})$, we generate a corresponding text representation by integrating layout specifications with content descriptions: $c^{\prime}=[c_{\text{layout}},c_{\text{content}}]$. Here, $c_{\text{layout}}$ encodes the spatial structure (e.g., a sequence arranged in $m\times n$ grids), while $c_{\text{content}}$ captures the visual content as well as the temporal relationships between frames.
@@
In this formulation, each frame $(I_{ij})_{t}$ interacts with others within the grid, enabling mutual influence. This interaction naturally enforces temporal consistency across the entire sequence.
@@ 2.3 Coarse-to-Fine Training
  • Coarse Phase: In the initial phase, we utilize large-scale Internet datasets, including WebVid, TikTok, and Objaverse, which are designed with uniform $c_{\text{layout}}$ specifications. Although the content descriptions ($c_{\text{content}}$) are automatically generated by GLM-4V-9B (Du et al., 2022) and may lack precise control details, the vast scale and diversity of this data—albeit at lower resolutions—provide a strong basis for developing robust spatial understanding and basic layout structures.
  • Fine Phase: Building on the foundational knowledge from the coarse phase, we transition to training with carefully curated, high-resolution samples. These samples are paired with detailed descriptions generated by GPT-4 (OpenAI, 2023), offering explicit spatial and temporal instructions. As shown in Figure 2, these high-quality captions facilitate fine-grained control over complex layout variations, enabling the model to handle intricate spatial and temporal dynamics effectively.
Loss Dynamic. Alongside the data curriculum, we gradually incorporate temporal supervision by employing a dynamic training objective:
@@
The base loss ensures accurate appearance generation for each grid position:
@@
The flow loss captures temporal dynamics through directional changes between adjacent positions:
$\Delta\epsilon^{ij}$ =$=\begin{cases}\epsilon^{ij}-\epsilon^{i,j-1}&\text{if }j>0\\ \epsilon^{i,0}-\epsilon^{i-1,n}&\text{if }j=0\end{cases}$ (5)
$\Delta\epsilon_{\theta}^{ij}$ =$=\begin{cases}\epsilon_{\theta}^{ij}-\epsilon_{\theta}^{i,j-1}&\text{if }j>0\\ \epsilon_{\theta}^{i,0}-\epsilon_{\theta}^{i-1,n}&\text{if }j=0\end{cases}$
@@ 2.4 Omni Inference
The generation process requires both flexibility and reference consistency. To achieve this balance, we introduce controlled grid noise injection instead of starting from pure noise:
@@
where $T$ denotes the starting time. This noise injection enables diverse generation while retaining the initialization structure.
To maintain reference consistency during generation, we employ a binary mask $M\in\{0,1\}^{m\times n}$:
@@
ensuring reference frames remain unchanged while allowing other regions to evolve.
The noise level $T$ plays a key role in balancing generation quality. A large $T$ leads to pure noise with poor reference consistency, while a small $T$ yields near-duplicates. Our experiments show $T\in[0.8,1.0]$ a good balance between diversity and fidelity.
@@ 3.1 Experimental Setup
We train our model separately for video generation and multi-view generation tasks, both following a two-stage strategy: (1) For coarse-level training, we combine video clips from WebVid (Bain et al., 2021), and TikTok (Jafarian & Park, 2022) arranged in 8×8 and 4×4 grid layouts for video generation, and 30K sequences from Objaverse (Deitke et al., 2023) in 4×6 grids for multi-view generation. Each sequence is paired with automated captions and GLM-generated annotations emphasizing spatial and temporal relationships, using the sequence’s inherent attributes (e.g., category labels) and visual content as queries. (2) For fine-grained control, we construct high-quality datasets of 1K sequences with structured annotations for each task. We first manually create exemplar annotations to establish a consistent format, then use these as few-shot examples for GPT-4o to generate precise control instructions while maintaining annotation consistency across the dataset.
@@
We implement GRID based on the FLUX-dev, initializing from its pretrained weights. For video generation training, we adopt LoRA with ranks of 16 and 256, training for 10K steps with batch size 4 across 8 A800 GPUs using AdamW optimizer (learning rate 1e-4). The temporal loss weight $\alpha$ starts from 0 and gradually increases to a maximum of 0.5. For multi-view generation, we train on 30K sequences for 1.5K steps using LoRA rank 256 and Ours-EF using LoRA rank 16. During inference, we use a guidance scale of 3.5 and sampling step of 20.
@@
We evaluate our model on three distinct generation tasks: (1) Text-to-video generation on UCF-101 dataset (Soomro et al., 2012), evaluated using FVD (Unterthiner et al., 2019) (I3D backbone) and IS (Xu et al., 2018). We evaluate both 16-frame and 64-frame generation settings; (2) Image-to-video generation on a randomly sampled subset of 100 TikTok videos, measured by FVD and CLIPimg score; (3) Multi-view generation on Objaverse, where we evaluate on 30 randomly selected objects with 24 frames per sequence at different viewpoints to assess 4D generation capabilities. We compute FVD, CLIP metrics, following (Liang et al., 2024).
@@ 3.2 Main Results
Table 1: Comprehensive Generation Results. Comparison across different generation tasks. Our model demonstrates superior efficiency while maintaining competitive quality across varying sequence lengths. Notably, while existing methods (AnimateDiffv3, VideoCrafter2) are limited to 16-frame generation, our approach scales efficiently to 64-frame sequences with only linear time increase. Underlined values indicate the best results among our variants, and bold values show the best across all methods. Test Time indicates the average sampling time per sequence.
Text-to-Video (16-frame) Text-to-Video (64-frame) Image-to-Video
[…]
AnimateDiffv3 464.1 35.24 12 - - - 250.9 0.9229 419M
VideoCrafter2 424.2 32.00 15 - - - - - 919M
OpenSora1.2 472.0 39.07 12 1000.5 37.11 66 - - 1.5B
CogVideo5b 301.1 36.27 48 740.1 34.82 132 122.5 0.9185 5B
Ours(Stage1) 482.1 32.46 7.2 1003.2 32.48 24 115.5 0.9598 160M
Ours(Stage1+2) 438.3 36.56 7.2 994.6 36.47 24 104.6 0.9695
Ours(Full) 418.9 37.34 7.2 721.6 36.63 24 93.7 0.9709
Table 2: Quantitative comparison of Multi-view Generation Results on Text-to-Multyview, Image-to-Multyview and Multyview interpolation Tasks. Time indicates the whole time cost during inference
Task Method CLIP-F↑ CLIP-O↑ FVD↓ Time↓
Text 2MV 4DFY 0.8092 0.6163 390.4 3h
Ours-EF 0.9060 0.6189 355.6 6m
Ours 0.9427 0.6247 324.3 6m
Image 2MV STAG4D 0.88 0.64 475.4 3h30m
Ours-EF 0.9392 0.6580 333.7 6m
Ours 0.9486 0.6554 350.6 6m
Interp olation Ours-EF 0.9356 0.7223 348.2 6m
Ours 0.9543 0.7415 295.1 6m
We compare our approach with several state-of-the-art methods from well-established video/multyview generation model series, all of which represent the current frontiers in their respective domains.
@@
As shown in Table 1, we evaluate our approach on both short (16-frame) and long (64-frame) generation tasks. For 64-frame generation, our full model pushes FVD to 721.6. For 16-frame, we significantly reduce computational costs, achieving 6.7× faster inference (7.2s vs 48s) compared to CogVideo. This efficiency advantage becomes more pronounced in 64-frame generation, where our approach maintains 5.5× faster inference speed. Our staged training demonstrates clear benefits: Stage1 (trained with coarse annotations and $\mathcal{L}_{base}$ only) achieves a FVD of 482.1, which reduces to 438.3 with fine-grained annotations, and further reduces to 418.9 with $\mathcal{L}_{flow}$.
@@
We evaluate on the TikTok dataset, which contains 100 high-quality short videos with diverse motion patterns after our processing and annotation process. For image-guided video synthesis, our method achieves breakthrough performance by reducing FVD to 93.7, a 23% improvement over the second-best result. The CLIPimg score also sees a substantial boost to 0.9709.
Multi-view Generation
We evaluate multi-view generation capabilities on selected Objaverse test set, which contains 30 3D objects from multiple viewpoints. As shown in Table 2, our method achieves transformative improvements in text-to-multi-view synthesis. We improve CLIP-F score by 13.4 points to 0.9427 and reduce FVD by 17% to 324.3, while dramatically cutting inference time from 3 hours to just 6 minutes compared to 4DFY (Bahmani et al., 2024). For image-to-multi-view tasks, we push the CLIP-F score up by 6.9 points to 0.9486 while achieving 35× faster inference compared to STAG4D (Zeng et al., 2024).

### v5

# Grid: Omni Visual Generation

@@ Abstract
Visual generation has witnessed remarkable progress in single-image tasks, yet extending these capabilities to temporal sequences remains challenging. Current approaches either build specialized video models from scratch with enormous computational costs or add separate motion modules to image generators, both requiring learning temporal dynamics anew. We observe that modern image generation models possess underutilized potential in handling structured layouts with implicit temporal understanding. Building on this insight, we introduce GRID, which reformulates temporal sequences as grid layouts, enabling holistic processing of visual sequences while leveraging existing model capabilities. Through a parallel flow-matching training strategy with coarse-to-fine scheduling, our approach achieves up to 67$\times$ faster inference speeds while using $<\frac{1}{1000}$ of the computational resources compared to specialized models. Extensive experiments demonstrate that GRID not only excels in temporal tasks from Text-to-Video to 3D Editing but also preserves strong performance in image generation, establishing itself as an efficient and versatile omni-solution for visual generation.
@@ 1 Introduction
Film strips demonstrate an elegant approach in visual arts: by arranging temporal sequences into structured grids, allowing time-based narratives to be displayed in layouts while maintaining their narrative coherence and visual connections. This organization does more than preserve chronological order - it enables efficient content manipulation, comparison, and editing. Drawing inspiration from this intuitive yet powerful organizational principle, we propose a fundamental question: Can we directly reframe various temporal visual generation tasks as how to layout, where key visual elements (such as multiple viewpoints or video frames) are treated as grid “layout”?
Figure 1: Different paradigms for temporal visual generation. (a) Motion-Scratch (e.g., SVD, AnimateDiff): learn temporal dynamics from scratch while reusing pretrained image models. (b) Full-Scratch (e.g., Sora): learn everything from scratch, requiring massive data and computational resources. (c) Zero-Scratch (GRID): reuse both spatial and temporal capabilities through grid-based reformulation, leveraging pretrained models’ inherent understanding.
To answer this, a natural starting point emerges from the recent breakthroughs in text-to-image generation. For single image generation, models like  [1, 2, 3] have demonstrated remarkable capabilities in understanding and generating complex spatial relationships. For temporal visual generation, current approaches typically follow two distinct paths: (a) building specialized video models from scratch (e.g., Sora), which requires learning both spatial and temporal relationships with prohibitive computational costs, (b) treating image generators as single-frame producers and mainly train additional motion modules - while this avoids learning spatial generation from scratch, it still requires learning temporal dynamics entirely anew.
Guided by our layout-centric perspective, we argue that the inherent capabilities of image generation models are significantly underestimated. Modern image models already possess implicit understanding of both spatial relationships and basic temporal coherence, suggesting we might not need to learn either aspect entirely from scratch. To validate this hypothesis, we first test the ability of current image generation models to handle grid-arranged layouts through simple prompting (Figure 4). Our experiments reveal that while these models show promising initial capabilities in understanding structured layouts, they still fall short in two fundamental aspects (detailed in Section A.1):
@@
Figure 2: Pipeline Overview. Left: GRID arranges videos into grid layouts, with text annotations combining layout format prefix and LLM-generated captions. The model is trained using LoRA fine-tuning on DIT blocks, incorporating both base loss and temporal loss to capture inter-frame relationships. Right: Grid-based reformulation naturally extends model’s built-in self-attention to include frame-wise self-attention, cross-frame attention, and text-to-frames cross-attention.
@@ 2.1 Grid Representation
Our approach demonstrates that thoughtful problem restructuring can be more effective than architectural modifications. By reorganizing the input space into a grid representation, standard text-to-image models can naturally handle temporal generation without architectural changes (see Appendix A.2 for detailed attention mechanism analysis). This grid-based design offers two key advantages: First, it enables parallel generation of all frames and eliminates the error accumulation problems common in autoregressive approaches [4]. Second, by leveraging the inherent consistency priors within pretrained image generation models, our approach effectively transfers their learned spatial consistency to temporal and multi-view coherence. This crucial advantage avoids the need for extensive pretraining on massive video datasets, as the grid representation naturally extends existing image-level understanding to sequence generation. Additionally, through flexible layout conditioning ($c_{\text{layout}}$), our model shows strong generalization capabilities beyond training constraints as shown in Appendix and Supplementary Material, suggesting a promising solution to the fixed-length limitations of existing methods. Additionally, our grid representation supports diverse input types, including multi-view images and multi-frame sequences, laying the foundation for a comprehensive omni-generation model that bridges image and video domains.
@@ 2.2 Parallel Flow Matching
To fully leverage the potential of our grid representation, we employ parallel flow matching [1] to ensure temporal coherence across consecutive grids. For each training sample $\mathbf{I}=(I_{ij})$, we generate a corresponding text representation by integrating layout specifications with content descriptions: $c^{\prime}=[c_{\text{layout}},c_{\text{content}}]$. Here, $c_{\text{layout}}$ encodes the spatial structure (e.g., a sequence arranged in $m\times n$ grids), while $c_{\text{content}}$ captures the visual content as well as the temporal relationships between frames.
@@
Each frame $(I_{ij})_{t}$ interacts with others within the grid, enabling mutual influence. This interaction naturally enforces temporal consistency across all sequences.
@@ 2.3 Coarse-to-Fine Training
  • Coarse Phase: In the initial phase, we utilize large-scale Internet datasets, including WebVid, TikTok, and Objaverse, which are designed with uniform $c_{\text{layout}}$ specifications. Although the content descriptions ($c_{\text{content}}$) are automatically generated by GLM-4V-9B [5] and may lack precise control details, the vast scale and diversity of this data—albeit at lower resolutions—provide a strong basis for developing robust spatial understanding and basic layout structures.
  • Fine Phase: Building on the foundational knowledge from the coarse phase, we transition to training with carefully curated, high-resolution samples. These samples are paired with detailed descriptions generated by GPT-4 [6], offering explicit spatial and temporal instructions. As shown in Figure 2, these high-quality captions facilitate fine-grained control over complex layout variations, enabling the model to handle intricate spatial and temporal dynamics effectively.
Loss Formulation. Our training objective combines appearance accuracy with temporal consistency through a weighted sum:
@@
The base loss ensures accurate noise prediction at each position using mean squared error:
@@
The flow loss enforces smooth temporal transitions by penalizing inconsistent changes between adjacent positions. For any position (i,j) in the grid, directional changes are:
$\Delta\epsilon^{ij}$ =$=\begin{cases}\epsilon^{ij}-\epsilon^{i,j-1}&\text{within row}\\ \epsilon^{i,0}-\epsilon^{i-1,n}&\text{across rows}\end{cases}$ (5)
Similarly for predicted values:
$\Delta\epsilon_{\theta}^{ij}$ =$=\begin{cases}\epsilon_{\theta}^{ij}-\epsilon_{\theta}^{i,j-1}&\text{within row}\\ \epsilon_{\theta}^{i,0}-\epsilon_{\theta}^{i-1,n}&\text{across rows}\end{cases}$ (6)
The flow loss then minimizes inconsistencies in these directional changes:
@@
Figure 3: Omni Inference Framework: By transforming temporal and view sequences into structured layout spaces, we enable a pure image-based model FLUX to tackle diverse video and multi-view tasks (text/image-to-video generation, video interpolation, and multi-view synthesis) through a unified pipeline without additional video-specific architectures.
@@ 2.4 Omni Inference
The generation process requires both flexibility and reference consistency. To achieve this, we introduce controlled grid noise injection instead of starting from pure noise:
@@
where $T$ denotes the time. This noise injection enables diverse generation while retaining the initialization structure. To maintain reference consistency during generation, we employ a binary mask $M\in\{0,1\}^{m\times n}$:
@@
ensuring reference frames remain unchanged while allowing other regions to evolve. The noise level $T$ plays a key role in balancing generation quality. A large $T$ leads to pure noise with poor reference consistency, while a small $T$ yields near-duplicates. Our experiments show $T\in[0.8,1.0]$ a good balance between diversity and fidelity.
@@ 3.1 Experimental Setup
We train our model separately for video generation and multi-view generation tasks, both following a two-stage strategy: (1) For coarse-level training, we combine video clips from WebVid [7], and TikTok [8] arranged in 8×8 and 4×4 grid layouts for video generation, and 30K sequences from Objaverse [9] in 4×6 grids for multi-view generation. Each sequence is paired with automated captions and GLM-generated annotations emphasizing spatial and temporal relationships, using the sequence’s inherent attributes (e.g., category labels) and visual content as queries. (2) For fine-grained control, we construct high-quality datasets of 1K sequences with structured annotations for each task. We first manually create exemplar annotations to establish a consistent format, then use these as few-shot examples for GPT-4o to generate precise control instructions while maintaining annotation consistency across the dataset.
@@
We implement GRID based on the FLUX-dev, initializing from its pretrained weights. For video generation training, we adopt LoRA with ranks of 16-256, training for 10K steps with batch size 4 across 8 A800 GPUs using AdamW optimizer (learning rate 1e-4). The temporal loss weight $\alpha$ starts from 0 and gradually increases to a maximum of 0.5. For multi-view generation, we train on 30K sequences for 1.5K steps using LoRA rank 256 and Ours-EF using LoRA rank 16. During inference, we use a guidance scale of 3.5 and sampling step of 20.
@@
We evaluate our model on three distinct generation tasks: (1) Text-to-video generation on UCF-101 dataset [10], evaluated using FVD [11] (I3D backbone) and IS [12]. We evaluate both 16-frame and 64-frame generation settings; (2) Image-to-video generation on a randomly sampled subset of 100 TikTok videos, measured by FVD and CLIPimg score; (3) Multi-view generation on Objaverse, where we evaluate on 30 randomly selected objects with 24 frames per sequence at different viewpoints to assess 4D generation capabilities. We compute FVD, CLIP metrics, following [13].
Table 1: Quantitative comparison of Multi-view Generation Results on Text-to-Multiview and Image-to-Multiview Tasks. Time indicates the whole time cost during inference in A800.
Text-to-Multiview (T2MV) Image-to-Multiview (I2MV)
Method CLIP-F↑ CLIP-O↑ FVD↓ Time↓ Method CLIP-F↑ CLIP-O↑ FVD↓ Time↓
Animate124 0.7889 0.6005 411.6 180m STAG4D 0.8803 0.6420 475.4 210m
4DFY 0.8092 0.6163 390.4 405m 4DGen 0.8724 0.6397 525.2 130m
Ours-EF 0.9060 0.6189 355.6 6m Ours-EF 0.9392 0.6580 333.7 6m
Ours 0.9427 0.6247 324.3 6m Ours 0.9486 0.6554 350.6 6m
Table 2: Comprehensive Generation Results. Our model achieves competitive quality with superior efficiency across tasks. While existing methods are limited to 16-frame generation, our approach efficiently scales to 64-frame sequences with linear time cost. Underlined and bold values indicate best results among our variants and all methods, respectively. Time shows average sampling time per sequence in A800 GPU. Para means training parameters.
[…]
AnimateDiffv3 464.1 35.24 12s - - - 250.9 0.9229 12s 419M
VideoCrafter2 424.2 32.00 15s - - - - - - 919M
OpenSora1.2 472.0 39.07 12s 1000.5 37.11 66s - - - 1.5B
Cosmos 399.7 35.54 275s - - - - - - 7B
CogVideo5b 410.5 38.8 13s 740.1 34.82 132s 122.5 0.9185 48s 5B
Ours(Stage1) 455.3 32.46 7.2s 1003.2 32.48 24s 115.5 0.9598 7.2s 160M
Ours(Stage1+2) 401.1 36.56 7.2s 994.6 36.47 24s 104.6 0.9695 7.2s
Ours(Full) 382.5 38.12 7.2s 721.6 36.63 24s 93.7 0.9709 7.2s
Table 3: Video Frame Interpolation Results on UCF101. We evaluate our full model following standard settings. All methods achieve comparable results, with our approach matching state-of-the-art EMA-VFI on PSNR.
Metrics ToFlow EMA-VFI UPR-Net VFIMamba Ours
[…]
@@ 3.2 Main Results
We compare our approach with state-of-the-art methods across key domains: multiview generation (Animate124 [14], 4DFY [15], STAG4D [16], 4DGen [17]), text-to-video and image-to-video generation (AnimateDiffv3 [18], OpenSora1.2 [19], Cosmos [20], CogVideo5b [21]), and video frame interpolation (EMA-VFI [22], UPR-NetV [23], FIMamba [24]).
Multi-view Generation
We evaluate on the Objaverse test set with 30 3D objects. As shown in Table 1, our method achieves state-of-the-art performance on both text-to-multiview and image-to-multiview tasks. For T2MV, we improve CLIP-F to 0.9427 and reduce FVD to 324.3, while achieving 67$\times$ faster inference (6m vs. 405m) compared to 4DFY. For I2MV, we achieve 0.9486 CLIP-F score with 35$\times$ speedup over STAG4D. Ours-EF (lora rank 16) also demonstrates strong performance-speed trade-off.
@@
As shown in Table 2, we achieve competitive FVD of 721.6 for 64-frame generation. For 16-frame generation, our method achieves 6.7$\times$ faster inference (7.2s vs 48s) compared to CogVideo, with the efficiency gap widening to 5.5$\times$ for 64-frame tasks. Our staged training shows clear progression: Stage1 achieves FVD 455.3, improving to 401.1 with fine-grained annotations, and further to 382.5 with $\mathcal{L}_{flow}$.
@@
We evaluate on the TikTok dataset containing 100 diverse short videos. Our method achieves breakthrough performance with FVD of 93.7 (23% improvement) and CLIPimg score of 0.9709. Notably, our approach requires only 160M parameters, compared to $>$400M for motion modeling or $>$1B for full generation in existing methods.
Video Frame Interpolation
We evaluate on the UCF101 dataset for video frame interpolation [22, 23, 24]. As shown in Table 3, our approach achieves state-of-the-art PSNR of 35.48, matching EMA-VFI. For SSIM, all methods perform comparably around 0.970.